# Bluesky and the AT Protocol: Usable Decentralized Social Media

### v1

@@ 3.3.2. The App View
To display this information in the user’s client app, the client queries the user’s own PDS, which then fetches the neccessary data from an App View. The App View is also responsible for enforcing moderation controls: for example, if one user has blocked another, and one of the users’ repositories contains a record of an interaction that should not have been allowed due to the block, then the App View drops that interaction so that nobody can see it in the client apps. This behavior is consistent with how blocking works on Twitter/X (X Help Center, [n. d.]), and it is also the reason why blocks are public records in Bluesky: every protocol-conforming App View needs to know who is blocking who in order to enforce the block (Bluesky Social PBC, 2023h; Newbold, 2023). If users are unhappy with the moderation rules applied by the App View operated by Bluesky Social PBC, it is always possible for third parties to operate alternative App Views that index the same firehose and present the data in a different way.
If the AT Protocol is used to implement another social mode besides microblogging, that application will most likely require an App View service of its own, which can be hosted by anyone. This service can then interpret and index the records in users’ repositories in whatever way is required for that application.
@@ 3.4. Labelers and Feed Generators
The AT Protocol seperates out the “opinionated” aspects of the system into separate services: labelers and feed generators. These services typically take the firehose as their input. Labelers produce a stream of judgements about content (e.g. “this post is spam”), whereas feed generators return a list of post IDs they have selected for inclusion in a custom feed, as described in Section 2.3. Users can choose in their client app which feeds and which labelers they want to use. The output of labelers is consumed by App Views or PDSes in order to apply content filtering (Bluesky Social PBC, 2023d). For a feed generator, an App View expands the post IDs into full posts before sending them to the client app of users who have subscribed to that feed.
@@
Feed generators can be implemented in code using our starter kit (Bluesky Social PBC, 2023b), or created with a third-party service such as Skyfeed (Bluesky Social PBC, 2023e).
@@ 3.5. User Identity
As explained in Section 2.2, user handles in Bluesky and atproto are DNS domain names. Any number of identity providers can coexist in the system: Bluesky Social PBC allows users to register subdomains of .bsky.social, but the indexing infrastructure does not treat users differently based on their handle.
We want a user to be able to change their handle without affecting their social graph. Therefore, when a record in user $A$’s repository indicates that $A$ is following $B$, that record must identify $B$ in a way that is more long-lived than specifying $B$’s handle. For this reason, every Bluesky/atproto account has an immutable, unique identifier: a decentralized ID or DID, which is a URI starting with the prefix did:. The record that $A$ follows $B$ then contains $B$’s DID. DIDs are a recent W3C standard (Sporny et al., 2022).
Moreover, we want a user to be able to migrate to a different PDS without changing either their DID or their handle. The DID specification provides a mechanism for resolving a DID into a DID document, a JSON document containing information about the user identified by that DID, as illustrated in Figure 4. In atproto, a DID document specifies (among other things) the handle of the user, the URL of their PDS, and the public key that is used to sign the Merkle tree root of their repository every time they add or delete a record. To change their handle or their PDS, the user needs to update their DID document to the new value.
For a user to successfully claim a particular handle, they must have a bidirectional link between their DID and their domain name handle, as shown in Figure 4:
@@ 3.5.1. Resolving DID documents
The W3C DID specification (Sporny et al., 2022) does not directly specify the mechanism for resolving a DID into a DID document. Rather, the first substring after did: in a DID indicates the DID method, and the specification of the DID method defines the protocol for obtaining the DID document. Hundreds of DID methods have been defined (Steele and Sporny, 2023), many of which are dependent on specific blockchains or other external systems. To avoid atproto implementations having to support so many resolution methods, our services currently only accept DIDs based on either did:web (defined by the the W3C Credentials Community Group (Gribneau et al., 2023)) or did:plc (defined by ourselves for atproto (Holmgren et al., 2023)). Support for more DID methods might be added in the future.
The did:web method is very simple: the part of the DID after did:web: is a domain name, and the DID document is resolved by making a HTTPS request to a /.well-known/ URL on that domain name (a path can optionally be included). The security of a did:web identity therefore assumes that the web hosting provider for that domain is trusted, and also relies on trusting the TLS certificate authorities that may authenticate the HTTPS request.
@@ 3.5.2. The did:plc DID method
The PLC directory server plays an authoritative role similar to the DNS root servers, but it is mostly untrusted because PLC DID documents are self-certifying. If the DID document has not changed since its initial creation, it is easy to verify that a DID has been correctly resolved to a DID document by recomputing its hash. To support changes to the DID document, the initial version of a user’s DID document contains a public key that is authorized to sign a new version of the DID document. Any new version of the DID document is only valid if it has been signed by the key in the previous version. The directory server returns all DID document versions for a given DID, allowing anybody to check the chain of signatures.
If the directory server were to be malicious, it would not be able to modify any DID documents – it could only omit valid DID document versions from its responses, or fail to respond at all. Moreover, if there were to be a fork in DID document history such that two correctly signed successor versions for some DID document exist, the directory server could choose which one of these forks to serve. To mitigate the risk of such attacks, we anticipate that a future version of the PLC directory will use techniques from certificate transparency (Laurie, 2014) to ensure that DID document updates form an append-only log.
@@ 3.5.3. Authentication
The Bluesky PDSes therefore hold these signing keys custodially on behalf of users, and users authenticate themselves to their home PDS via username and password. This provides a familiar user experience to users, and enables standard features such as password reset by email. The AT Protocol does not make any assumptions about how PDSes authenticate their users, and other PDS operators are free to use different authentication methods.
@@ 4. Related Work
Several other decentralized social networks are in development. We believe that there is no single optimal design: different systems make different trade-offs, and are therefore suitable for different purposes. Bluesky and the AT Protocol aim to provide a good user experience by making moderation a first-class concern, by having clients that are lightweight and fast, and by providing a global view over the whole network. For example, conversation threads include all replies (unless removed by moderation), regardless of the server on which they were posted. To achieve this goal we rely on an indexing infrastructure that is more centralized than some other designs. However, we emphasize that there can be multiple competing indexers, and third-party client apps are free to show data from whichever indexers they wish.
@@
Many decentralized social networking projects have ideas in common. For example, the idea of using DNS domain names as usernames also appears in Nostr (fiatjaf and Dilger, 2021). An atproto PDS has similarities to Git repository hosting (e.g. GitHub/Gitlab) or a Solid pod (Solid Project, [n. d.]). There are also federated chat systems such as Matrix (Matrix.org Foundation, [n. d.]), IRC (Oikarinen and Reed, 1993), and XMPP (Saint-Andre, 2011), but we focus on systems that provide a Twitter-like model where users follow each other.
@@ 4.1. Scuttlebutt
Secure Scuttlebutt (SSB) is a peer-to-peer social networking protocol (Scu, [n. d.]); Manyverse (The Manyverse Authors, [n. d.]c) and Planetary (Verse Communications Inc, [n. d.]) are social applications built upon the SSB protocol. It optionally uses relay servers called pubs to store messages from peers that are offline, and to enable user discovery. The client software downloads the feeds from accounts that the user is explicitly following, and from accounts followed by followed accounts (up to three hops by default). This can require significant amounts of storage and bandwidth on the client.
Any messages from users outside of the third-degree network are not shown, which effectively limits the set of people who can mention or reply to a user to the third-degree network. This deliberate design decision is intended to reduce moderation problems by prioritizing conversation between people who already know each other (Staltz, 2023b). In contrast, Bluesky/atproto are designed to allow anybody to talk to anybody else. This requires more explicit moderation to manage unwanted content, but we believe it also enables serendipity and is a prerequisite for any “digital town square”.
Since SSB is built upon append-only logs and gossip replication, it is not possible to delete content once it has been posted (The Manyverse Authors, [n. d.]b). User identity is tied to a cryptographic key on the user’s device, requiring manual key management for moving to another device. Posting from multiple devices is not possible, as sharing the same key between devices can make an account unrecoverable (The Manyverse Authors, [n. d.]a). A work-in-progress successor protocol to SSB, called PPPPP, is designed to address these issues (Staltz, 2023a).
@@ 4.2. Nostr
Nostr also began as a revision of SSB, replacing the append-only logs with individual signed messages (Henshaw-Plath, 2023). It leans more heavily on relay servers instead of peer-to-peer communication: clients publish and fetch messages on relays of their choice, and there is no federation among relays (fiatjaf, [n. d.]). The protocol is deliberately simple, and it prioritizes censorship resistance over moderation: relays can block users, but users can always move to a new relay, and use multiple relays at the same time. Communication (e.g. reply threads) is only possible between users who have at least one relay in common. Although some services index the whole Nostr network, these indexes are not used for user-to-user interaction. As a result, it is unpredictable who will see which message. The creator of Nostr writes: “there are no guarantees of anything […] to use Nostr one must embrace the inherent chaos” (fiatjaf, 2023). Key management is manual in Nostr, and facilities for key rotation are still under discussion (Ball et al., 2022).
@@ 4.3. Farcaster and blockchain-based systems
Farcaster (Warpcast, [n. d.]) has some architectural similarities to Bluesky/atproto, although it was developed independently. It has storage servers called hubs, which store the state of the entire network similarly to an atproto Relay, and it has a concept of hosted app servers that are similar to our App View (Srinivasan, [n. d.]). Farcaster user IDs are similar to our DIDs, and they are mapped to public keys using a smart contract on the Ethereum Optimism blockchain that is functionally similar to our PLC directory. Usernames can be either ENS names (ENS Labs Limited, [n. d.]), or names within an off-chain namespace managed centrally by Farcaster, similarly to .bsky.social subdomains in Bluesky (Farcaster Team, [n. d.]).
@@
The Lens protocol (Lens Protocol, [n. d.]) is more strongly blockchain-based than Farcaster: it even stores high-volume user actions such as posts and follows on Polygon, a proof-of-stake blockchain. DSNP takes a similar approach (Project Liberty, 2020). Placing high-volume events directly on a blockchain incurs orders of magnitude higher per-user costs than atproto, and is likely to run into scalability limits as the number of users grows. Lens is adopting a layer-3 blockchain that provides better scalability and lower cost (Lens Protocol, 2023), but weaker security properties. Linking social accounts to cryptocurrency wallets and NFTs enables users to monetize their content, but this is not a goal of atproto.
@@ 4.4. ActivityPub and Mastodon
ActivityPub (Lemmer-Webber et al., 2018) is a W3C standard for social networking, and Mastodon (Mastodon gGmbH, [n. d.]) is its most popular implementation. We have highlighted aspects of their design in Sections 1, 2.3, and 3.2. Mastodon gives a lot of power to server administrators: for example, a server admin can choose to block another server, preventing all communication between users on those servers. There is a degree of lock-in to a server because moving to another server is intrusive: the username changes, moving posts to the new server currently requires an experimental command-line tool (SilverWolf32 et al., 2019; Tokyo Outsider, 2023), and other users’ replies to those posts are lost. These risks can be mitigated by self-hosting; managed providers exist (Grow your own services, [n. d.]), but they still require some expertise and cost money. The AT Protocol separates the roles of moderation and hosting, and aims to make it easier to change providers.
When user $A$ follows user $B$, AA’s server asks $B$’s server to send it notifications of $B$’s future posts via ActivityPub. This architecture has the advantage of not requiring a whole-network index. However, replies to a post notify the server of the original poster, but not necessarily every server that has a copy of the original post, leading to inconsistent reply threads on different servers. Notifications can be forwarded, but in the limit this leads to each server having a copy of the whole network, which would make it expensive to run a server. Viral posts can generate a lot of inbound requests to a server from people liking, replying, and boosting (reposting). The Bluesky indexing infrastructure is also fairly expensive, but a PDS is cheap to run. Since users can choose their moderation preferences independently from their indexing provider (App View), we believe that the ecosystem can be healthy with a small number of indexing providers.
@@ 5. Conclusions
Bluesky and the AT Protocol are a new approach to social media. While some decentralized systems prioritize censorship resistance, we believe that a good user experience requires explicitly addressing problematic content such as harassment and misinformation. We therefore make moderation a first-class concern that is handled separately from infrastructure hosting, and we provide strong mechanisms for users to control the content they see.
Our open architecture allows a pluralistic system in which there is no global consensus on what content is acceptable. A user on a self-hosted or loosely moderated PDS may post controversial content, but they are not entitled to the attention of others: App Views may choose not to index the content, and clients may ignore it depending on the user’s moderation settings. This philosophy is sometimes described as “free speech, but not free reach” (DiResta, 2018).
The AT Protocol provides cryptographically authenticated data, but our implementation pairs it with custodial key management to provide a familiar user experience. The system is open to third-party clients and alternative PDS hosts, and anybody can index the network using real-time streams for low-latency updates. This reduces the dependency on any one provider, since every part of the system can be run by multiple competing providers, and users can switch providers with minimal friction (in particular, without changing username, and without losing any of their content or social graph). The AT Protocol is also extensible to other social modes besides microblogging.
Acknowledgements.
Martin Kleppmann is supported by the Volkswagen Foundation and crowdfunding supporters including Mintter and SoftwareMill.

### v2

@@ 3.3.2. The App View
To display this information in the user’s client app, the client queries the user’s own PDS, which then fetches the necessary data from an App View. The App View is also responsible for enforcing moderation controls: for example, if one user has blocked another, and one of the users’ repositories contains a record of an interaction that should not have been allowed due to the block, then the App View drops that interaction so that nobody can see it in the client apps. This behavior is consistent with how blocking works on Twitter/X (X Help Center, [n. d.]), and it is also the reason why blocks are public records in Bluesky: every protocol-conforming App View needs to know who is blocking who in order to enforce the block (Bluesky Social PBC, 2023h; Newbold, 2023). If users are unhappy with the moderation rules applied by the App View operated by Bluesky Social PBC, it is always possible for third parties to operate alternative App Views that index the same firehose and present the data in a different way.
Other social modes built upon atproto require their own App View services to interpret and index the records in users’ repositories as required. The PDSes and relay can be reused.
@@ 3.4. Labelers and Feed Generators
The AT Protocol separates out the “opinionated” aspects of the system into separate services: labelers and feed generators. These services typically take the firehose as their input. Labelers produce a stream of judgements about content (e.g. “this post is spam”), whereas feed generators return a list of post IDs they have selected for inclusion in a custom feed. Users can choose in their client app which feeds and which labelers they want to use. The output of labelers is consumed by App Views or PDSes in order to apply content filtering (Bluesky Social PBC, 2023d). For a feed generator, an App View expands the post IDs into full posts before sending them to the client app of users who have subscribed to that feed.
@@
Feed generators can be implemented using our starter kit (Bluesky Social PBC, 2023b), or using a third-party service such as Skyfeed (Bluesky Social PBC, 2023e).
@@ 3.5. User Identity
The DNS-based user identity in atproto is decoupled from hosting and indexing. Anybody owning a domain name can issue subdomains as atproto handles, and this does not require running a PDS or any other service besides DNS.
Handles can change, but every Bluesky/atproto account also has an immutable, unique identifier: a decentralized ID or DID, which is a URI starting with the prefix “did:”. For example, when a record in user $A$’s repository indicates that $A$ is following $B$, that record contains $B$’s DID; this allows a user to change their handle without affecting their social graph. DIDs are a recent W3C standard (Sporny et al., 2022).
Moreover, we want a user to be able to migrate to a different PDS without changing either their DID or their handle. DIDs provide a mechanism for resolving a DID into a DID document, a JSON document containing information about the user identified by that DID, as illustrated in Figure 2. In atproto, a DID document specifies the handle of the user, the URL of their PDS, and the public key that is used to sign the Merkle tree root of their repository every time they add or delete a record. To change their handle or their PDS, the user needs to update their DID document to the new value.
To prove ownership of a handle, the user must have a bidirectional link between their DID and their domain name handle, as shown in Figure 2:
@@
The App View periodically checks these bidirectional links, and invalidates the handle if either is broken. Provided that the App View is honest and takes measures to protect against DNS poisoning attacks (perhaps using DNSSEC when available), this approach prevents users from impersonating domains that they do not own.
@@ 3.5.1. Resolving DID documents
The DID specification (Sporny et al., 2022) does not directly specify the mechanism for resolving a DID into a DID document. Rather, the first substring after did: in a DID indicates the DID method, and the specification of the DID method defines the protocol for obtaining the DID document. Hundreds of DID methods have been defined (Steele and Sporny, 2023), many of which are dependent on specific blockchains or other external systems. To avoid atproto implementations having to support so many resolution methods, our services currently only accept DIDs based on either did:web (defined by the the W3C Credentials Community Group (Gribneau et al., 2023)) or did:plc (defined by ourselves for atproto (Holmgren et al., 2023)). Support for more DID methods might be added in the future.
The did:web method is very simple: the part of the DID after did:web: is a domain name, and the DID document is resolved by making a HTTPS request to a /.well-known/ URL on that domain name. The security of a did:web identity therefore assumes that the web hosting provider for that domain is trusted, and also relies on trusting the TLS certificate authorities that may authenticate the HTTPS request.
@@ 3.5.2. The did:plc DID method
The PLC directory server plays an authoritative role similar to the DNS root servers, but it is mostly untrusted because PLC DID documents are self-certifying. If the DID document has not changed since its initial creation, it is easy to verify that a DID has been correctly resolved to a DID document by recomputing its hash. To support changes to the DID document, the initial version of a user’s DID document contains a public key that is authorized to sign a new version of the DID document. Any new version of the DID document is only valid if it has been signed by the key in the previous version. The directory returns all DID document versions for a given DID, allowing anybody to check the chain of signatures.
If the directory server were to be malicious, it would not be able to modify any DID documents – it could only omit valid DID document versions from its responses, or fail to respond at all. Moreover, if there were to be a fork in DID document history such that two correctly signed successor versions for some DID document exist, the directory server could choose which one of these forks to serve. To mitigate such attacks, we anticipate that a future version of the PLC directory will use an append-only transparency log similar to certificate transparency (Laurie, 2014).
@@ 3.5.3. Authentication
The Bluesky PDSes therefore hold these signing keys custodially on behalf of users, and users log in to their home PDS via username and password. This provides a familiar user experience to users, and enables standard features such as password reset by email. The AT Protocol does not make any assumptions about how PDSes authenticate their users; other PDS operators are free to use different methods, including user-managed keys.
@@ 4. Related Work
Several other decentralized social networks are in development. We believe that there is no single optimal design: different systems make different trade-offs, and are therefore suitable for different purposes. Bluesky and the AT Protocol aim to provide a good user experience by providing a global view over the whole network, making moderation a first-class concern, and having clients that are lightweight and easy to use. For example, conversation threads include all replies (unless removed by moderation), regardless of the server on which they were posted. To achieve this goal we rely on an indexing infrastructure that is more centralized than some other designs. However, we emphasize that there can be multiple competing indexers, and third-party client apps are free to show data from whichever indexers they wish.
@@
Many decentralized social networking projects have ideas in common. For example, the idea of using DNS domain names as usernames also appears in Nostr (fiatjaf and Dilger, 2021). An atproto PDS has similarities to Git repository hosting (e.g. GitHub/Gitlab) or a Solid pod (Solid Project, 2024).
@@ 4.1. Scuttlebutt
Secure Scuttlebutt (SSB) is a peer-to-peer social networking protocol (Scu, 2023); Manyverse (The Manyverse Authors, 2024) is a social application built upon the SSB protocol. It optionally uses relay servers called pubs to store messages from peers that are offline, and to enable user discovery. The client software downloads the feeds from accounts that the user is explicitly following, and from accounts followed by followed accounts (up to three hops by default). This can require significant amounts of storage and bandwidth on the client.
Any messages from users outside of the third-degree network are not shown, which effectively limits the set of people who can mention or reply to a user to the third-degree network. This deliberate design decision is intended to reduce moderation problems by prioritizing conversation between people who already know each other (Staltz, 2023). In contrast, Bluesky/atproto are designed to allow anybody to talk to anybody else. This requires more explicit moderation to manage unwanted content, but we believe it also enables serendipity and is a prerequisite for any “digital town square”.
Since SSB is built upon append-only logs and gossip replication, it is not possible to delete content once it has been posted (The Manyverse Authors, 2021). User identity is tied to a cryptographic key on the user’s device, requiring manual key management for moving to another device. Posting from multiple devices is not possible, as sharing the same key between devices can make an account unrecoverable (The Manyverse Authors, 2022). A successor protocol to SSB, called PZP, is designed to address these issues (Karlsson, 2024).
@@ 4.2. Nostr
Nostr also began as a revision of SSB, replacing the append-only logs with individual signed messages (Henshaw-Plath, 2023). It leans more heavily on relay servers instead of peer-to-peer communication: clients publish and fetch messages on relays of their choice, and there is no federation among relays (fiatjaf, 2024). The protocol is deliberately simple, and it prioritizes censorship resistance over moderation: relays can block users, but users can always move to a new relay, and use multiple relays at the same time. Communication (e.g. reply threads) is only possible between users who have at least one relay in common. Although some services index the whole Nostr network, these indexes are not used for user-to-user interaction. As a result, it is unpredictable who will see which message. The creator of Nostr writes: “there are no guarantees of anything […] to use Nostr one must embrace the inherent chaos” (fiatjaf, 2023). Key management is manual in Nostr, and facilities for key rotation are still under discussion (Ball et al., 2022).
@@ 4.3. Farcaster and blockchain-based systems
Farcaster (Warpcast, 2024) has some architectural similarities to Bluesky/atproto, although it was developed independently. It has storage servers called hubs, which store the state of the entire network similarly to an atproto Relay, and it has a concept of hosted app servers that are similar to our App View (Srinivasan, 2023). Farcaster user IDs are similar to our DIDs, and they are mapped to public keys using a smart contract on the Ethereum Optimism blockchain that is functionally similar to our PLC directory. Usernames can be either ENS names (ENS Labs Limited, 2024), or names within an off-chain namespace managed centrally by Farcaster, similarly to .bsky.social subdomains in Bluesky (Farcaster Team, 2024).
@@
The Lens protocol (Lens Protocol, 2024) is more strongly blockchain-based than Farcaster: it even stores high-volume user actions such as posts and follows on Polygon, a proof-of-stake blockchain. DSNP takes a similar approach (Project Liberty, 2020). Placing high-volume events directly on a blockchain incurs orders of magnitude higher per-user costs than atproto, and is likely to run into scalability limits as the number of users grows. Lens is adopting a layer-3 blockchain that provides better scalability and lower cost (Lens Protocol, 2023), but weaker security properties. Linking social accounts to cryptocurrency wallets and NFTs enables users to monetize their content, but this is not a goal of atproto.
@@ 4.4. ActivityPub and Mastodon
ActivityPub (Lemmer-Webber et al., 2018) is a W3C standard for social networking, and Mastodon (Mastodon gGmbH, 2024) is its most popular implementation. Mastodon gives a lot of power to server administrators: for example, a server admin can choose to block another server, preventing all communication between users on those servers. There is a degree of lock-in to a server because moving to another server is intrusive: the username changes, moving posts to the new server currently requires an experimental command-line tool (SilverWolf32 et al., 2019; Tokyo Outsider, 2023), and other users’ replies to those posts are lost. If the old server is not reachable – for example, because its admin shut it down without warning or because its domain was seized (Erin, 2024) – the user’s social graph is lost. These risks can be mitigated by self-hosting; managed providers exist (Grow your own services, 2024), but they still require some expertise and cost money. The AT Protocol separates the roles of moderation and hosting, and aims to make it easier to change providers without losing any data.
When user $A$ follows user $B$, AA’s server asks $B$’s server to send it notifications of $B$’s future posts via ActivityPub. This architecture has the advantage of not requiring a whole-network index. However, replies to a post notify the server of the original poster, but not necessarily every server that has a copy of the original post, leading to inconsistent reply threads on different servers. Notifications can be forwarded, but in the limit this leads to each server having a copy of the whole network, which would make it expensive to run a server. Viral posts can generate a lot of inbound requests to a server from people liking, replying, and boosting (reposting). In comparison, the Bluesky indexing infrastructure is also fairly expensive, but a PDS is cheap to run. Since users can choose their moderation preferences independently from their indexing provider (App View), we believe that the ecosystem can be healthy with a small number of indexing providers.
@@ 5. Conclusions
Bluesky and the AT Protocol are a new approach to social media. Their architecture is based on the principle that every part of the system can be run by multiple competing providers, and users can switch providers with minimal friction (in particular, without changing username, and without losing any of their content or social graph). For example, anyone can write a client, host a PDS, index the network, or provide moderation services, and all of these services interoperate. Even though the majority of Bluesky services are currently operated by a single company, we nevertheless consider the system to be decentralized because it provides credible exit (Newbold, 2024b): if Bluesky Social PBC goes out of business or loses users’ trust, other providers can step in to provide an equivalent service using the same dataset and the same protocols.
While some decentralized systems prioritize censorship resistance, we believe that a good user experience requires explicitly addressing problematic content such as harassment and misinformation. We therefore make moderation a first-class concern that is handled separately from infrastructure hosting, and we provide strong mechanisms for users to control the content they see. Our open architecture allows a pluralistic system in which different users may choose different providers that uphold different values, while still allowing them to communicate and interoperate.